# Supervisory Measurement-Guided Noise Covariance Estimation

### v1

@@ I INTRODUCTION
However, obtaining precise knowledge of noise covariance is challenging [1], as sensor characteristics vary with environmental conditions and often require recalibration. Moreover, front-end preprocessing, such as feature extraction, alters raw measurements and obscures their true statistical properties [2]. Noise covariance estimation has consequently received significant research attention. Existing approaches span from MLE and MAP formulations that explicitly optimize noise covariances within probabilistic models [2, 3], to performance-driven methods that tune parameters by minimizing trajectory tracking error [4, 5]. Both gradient-based methods leveraging analytical derivative solutions [6] and derivative-free approaches that avoid explicit gradient computation [7, 8, 9] have been explored.
@@ II Related Works
1) MAP/MLE methods: Noise covariance optimization via MLE/MAP formulations has been extensively studied. While derivative-free methods like expectation maximization jointly optimize parameters and states under likelihood objectives [7, 8], they require full posterior approximation, while point estimation is computationally preferable in practice. Gradient-based methods for parameter estimation in Kalman filtering minimize the negative log-likelihood, also known as the energy function [6], using only the observations directly employed by the filter. The required gradients can be obtained either through forward sensitivity equations [3] or via backpropagation [10]. In [2], states and noise covariances are jointly optimized under a MAP formulation by exploiting the convexity of the joint estimation problem and using an alternating optimization scheme, with most of the computational cost incurred in solving for the states.
2) Performance-driven methods: Several existing works estimate noise parameters under the supervision of state-estimation performance. The method in [11] optimizes innovation-based objectives, using measurement residuals in the outer loop and state estimation in the inner loop. The work in [12] proposes an adaptive disturbance estimation for a moving horizon estimator via bilevel optimization with fast derivative computation, while it requires a reference trajectory. The framework in [4] formulates covariance learning as constrained bilevel optimization, which requires ground-truth supervision and uses numerical gradients. Parameters can be optimized to fit the current task using these approaches. However, without probabilistic modeling, the resulting parameters may deviate from the true underlying values.
@@ III Preliminary on Robotic Lie Groups
This section presents the preliminaries on matrix Lie groups used in the derivation of this work, following [13].
@@
Let $\mathcal{M}$ be the smooth manifold of a Lie group. For any $X_{1},X_{2}\in\mathcal{M}$ and tangent vector $\xi\in\mathbb{R}^{\dim\mathcal{M}}$, we define the plus and minus operators, $X_{1}\boxplus\xi$ and $X_{1}\boxminus X_{2}$, which map between the manifold and its tangent space. These operators admit left and right versions; the choice is application-dependent and left implicit. For instance, in the left case, $X_{1}\boxplus\xi=\exp(\xi){X}_{1},X_{1}\boxminus X_{2}=\log(X_{1}X_{2}^{-1}).$ On product manifolds, the operators act componentwise [13]. The uncertainty of $X\in\mathcal{M}$ is modeled in the tangent space by a Gaussian distribution [14]. Let $\xi\sim\mathcal{N}(0,\Sigma)$, the induced distribution is
@@ IV Problem Statement
where $X_{k}\in\mathcal{X}$ denotes the robot’s state at time step $k$ on the Lie group $\mathcal{X}$, $u_{k}\in\mathcal{U}$ is the control input, and $w_{k}$ is the process noise. Conventionally, introspective sensor measurements (e.g., from IMUs) are used as inputs to the motion model through preintegration, which conflates uncertainty in measurements with the process noise $w_{k}$ [15].
The motion of a robot can be captured using extrospective sensors. In this paper, we categorize their measurements into two types [16]. Unary measurements, including those from GPS receivers, correspond to a single motion state and provide information in a fixed reference frame, whereas binary measurements, including LiDAR or RGB-D point clouds from iterative closest point (ICP), relate two consecutive motion states. At this stage, we do not constrain the sensor types to specific implementations to maintain theoretical completeness. The probabilistic models for these two measurement classes are
@@
As discussed in Section I, accurate knowledge of process and measurement noise covariances is crucial for odometry but remains difficult to determine in practice. Consequently, we treat these covariances as unknown quantities and introduce a parameterization via a variable $\theta$. This parameterization may take different forms, from simple vectorization of matrix entries (with symmetry, positive definiteness, and other constraints enforced) to more expressive representations based on neural networks. While some prior works [4, 5] leverage ground-truth robot poses as supervisory signals to learn the noise parameters, in this paper, we propose a general approach. We introduce a new class of measurements—termed supervisory measurements—that directly inform the robot’s relative poses over extended trajectories. These measurements provide supervision for estimating $\theta$. Formally, the supervisory measurement model is
@@ V Methodology
The assumption is standard in SLAM literature [2, 17], which enables tractable theoretical analysis while remaining practically justifiable for most real-world applications.
@@ V-B State Filter
where $F_{k-1}=\bigl[\begin{smallmatrix}\Phi_{k-1}&\bm{0}\\ \bm{0}&I\end{smallmatrix}\bigr]$, $Q_{0,k}=\bigl[\begin{smallmatrix}Q_{0}&\bm{0}\\ \bm{0}&\bm{0}\end{smallmatrix}\bigr]$, with $\Phi_{k-1}\triangleq\bigl[\begin{smallmatrix}\bm{0}&I\\ \bm{0}&\phi_{k-1}\end{smallmatrix}\bigr]$, $Q_{0}\triangleq\bigl[\begin{smallmatrix}\bm{0}&\bm{0}\\ \bm{0}&Q\end{smallmatrix}\bigr]$. Note that $\phi_{k-1}$ follows [18] and owing to the log-linear property of group-affine dynamics, $\phi_{k-1}$ is independent of ${X}_{k-1}$.
@@
By substituting the Gaussian predictive distribution from Subsection V-B into (9) and discarding additive constants, the odometry loss takes the closed-form expression [6]
@@
Evaluating the gradient of $\ell^{o}(\theta)$ with respect to $\theta$ requires the derivatives of $S_{k}$ and $r_{k}$, with respect to $\theta$, which in turn necessitates a sensitivity analysis of the state filter. Following [6], applying a term-wise differential to the state filter (11)–(17), which we refer to as the derivative filter.
@@ V-D Methodology Summary and More Discussions
The overall execution of the noise–covariance estimation procedure is summarized in Algorithm 1. At the upper level, the problem is posed as a constrained optimization, where the operation $\mathrm{Projection}_{\Theta}$ projects the estimated parameters onto the feasible set $\Theta$. To improve numerical stability, feasibility constraints can incorporate bounds on the eigenvalues or the condition number of the covariance matrix [4]. In this work, the step size $\eta^{(i)}$ is obtained using the Armijo line search rule. More generally, once the objective function and its gradient are available, a wide range of gradient-based optimization algorithms can be employed.
@@
Multiple parameterizations are possible for noise covariance matrices. A widely adopted approach is Cholesky decomposition [10], which inherently guarantees positive definiteness, while the diagonal form [4] is a special case. Our framework is agnostic to the specific choice of parameterization and can accommodate more expressive representations, including differentiable neural network layers [12].
@@
The proposed method is not limited to the InEKF. For a general EKF, the required derivatives can be obtained via sensitivity equations. We focus on the InEKF and affine group systems in this work, as they are prevalent in practice and possess favorable linearization properties [18] that simplify the derivative–filter design.
@@
Finally, we briefly relate our approach to prior work. Noise covariance estimation is often posed as alternating optimization over states and covariances, using either a shared or separate objective at two levels. We start with the covariance-level objective $L(\theta)$, which is the log-likelihood of the parameters of interest conditioned on all measurements, as derived from a Bayesian framework. Serval prior works [4, 5, 12] tune hyperparameters, including noise covariances, by minimizing state-estimation losses against ground-truth poses, $L(\hat{\bm{X}}(\theta))$. Although effective in reducing trajectory error, the resulting covariances may not reflect the true noise statistics well. To remove reliance on ground truth, [19] introduced the approximate posterior error, which in the EKF reduces to the trace of the estimation error covariance, i.e., $L(\hat{\bm{P}}(\theta))$. Innovation-based methods learn covariances from measurement residuals, leading to $L(\hat{\bm{X}}(\theta))$. The traditional energy minimization approach [3, 6] corresponds to the loss $\ell^{o}(\theta)$, whereas our formulation incorporates the prevalent supervisory measurements in SLAM as an additional source of information to reinforce estimation performance. More recently, [2] proposed joint MAP estimation of states and noise covariances with closed-form covariance updates, but this framework is incompatible with filter-based state estimation and entails substantial cost due to repeated full-state optimization refinement.
@@ VI-A Baselines
    Approximate posterior error (APE) [9].
@@
    Mean squared error of state estimate (MSE) [4].
@@
    Average innovartion (Innov.) [11].
@@
The implementations of gradient descent optimization in [4, 11] follow the original papers, which rely on numerical gradients. The approach of [9] uses Bayesian optimization for general hyperparameters, whereas we focus on a differentiable noise parameter and apply gradient descent.
@@ VI-B Simulation
We first evaluate the algorithm in simulation. The state $X_{k}\in SE_{2}(3)$ as in (1) is estimated using a left-invariant extended Kalman filter (LInEKF). The system kinematics is given by the IMU, where in LInEKF the process noise covariance is $\mathrm{blkdiag}(Q_{g},\bm{0},Q_{a})$ [20], with gyroscope noise $n_{g}\sim\mathcal{N}(\bm{0},Q_{g})$ and accelerometer noise $n_{a}\sim\mathcal{N}(\bm{0},Q_{a})$. The unary measurements given by GPS and binary measurements given by visual odometry (VO) are specified as
@@
A Monte Carlo experiment is conducted to evaluate the impact of supervisory measurements by comparing $\ell^{o}$ and $\mathcal{L}$ as loss functions. Measurement noises are parameterized by a base level $\beta$ and a scaling factor $\alpha$, with $\theta_{j}=(\alpha\beta_{j})^{2}$, where $\beta_{j}$ is sampled from $1,2]\times 10^{-4}$ for $Q$, $2,10]\times 10^{-3}$ for $\Sigma^{u}$, and $1,6]\times 10^{-4}$ for $\Sigma^{b}$. All algorithms are initialized identically and run for 20 iterations under the same optimizer settings. The average 2-Wasserstein error [2] is also reported, measuring the discrepancy between estimated and true covariances, with smaller values indicating more accurate estimation. As shown in Figure.5, $\mathcal{L}$ yields consistently lower MSE than $\ell^{o}$ across noise levels and, under large noise, achieves smaller Wasserstein error, indicating more accurate covariance estimation. This demonstrates the additional benefits of incorporating supervisory measurements.
@@ VI-C Dataset Experiment
The garage dataset [21] provides real-world binary measurements and loop closures for evaluation. It features calibration segments with abundant loops and validation sections with open-loop trajectories. Lacking unary measurements, we use only IMU and binary data. Following convention, we treat the outlier-free estimation using all measurements as groundtruth. Noisy IMU measurements are then generated from this trajectory using known noise parameters, enabling convenient evaluation of the algorithm. As outlier rejection is not the focus of this work, we replace the false positive loop closures with measurements assigned a noise standard deviation of $0.05$. In the loop- case, 136 loop pairs are used, while in the loop+ case, 234 pairs are included.
@@ VII Conclusion and Future Work
Several aspects of the framework warrant further exploration. We plan to apply the algorithm to real-world SLAM systems such as [17]. Besides, we will explore deployment-oriented parameterizations, for instance, differentiable neural embeddings that adapt measurement noises to context. Moreover, the framework’s flexibility also enables tuning of broader hyperparameters, which remains to be investigated.

### v2

@@ I INTRODUCTION
However, obtaining precise knowledge of noise covariance is challenging [ebadi2023present], as sensor characteristics vary with environmental conditions and often require recalibration. Moreover, front-end preprocessing, such as feature extraction, alters raw measurements and obscures their true statistical properties [khosoussi2025joint]. Noise covariance estimation has consequently received significant research attention. Existing approaches span from MLE and MAP formulations that explicitly optimize noise covariances within probabilistic models [khosoussi2025joint, tsyganova2017svd], to performance-driven methods that tune parameters by minimizing trajectory tracking error [qadri2024learning, liu2025debiasing]. Both gradient-based methods leveraging analytical derivative solutions [sarkka2023bayesian] and derivative-free approaches that avoid explicit gradient computation [9357964, burnett2021radar, hu2017efficient] have been explored.
@@ II Related Works
1) MAP/MLE methods: Noise covariance optimization via MLE/MAP formulations has been extensively studied. While derivative-free methods like expectation maximization jointly optimize parameters and states under likelihood objectives [9357964, burnett2021radar], they require full posterior approximation, while point estimation is computationally preferable in practice. Gradient-based methods for parameter estimation in Kalman filtering minimize the negative log-likelihood, also known as the energy function [sarkka2023bayesian], using only the observations directly employed by the filter. The required gradients can be obtained either through forward sensitivity equations [tsyganova2017svd] or via backpropagation [parellier2023speeding]. In [khosoussi2025joint], states and noise covariances are jointly optimized under a MAP formulation by exploiting the convexity of the joint estimation problem and using an alternating optimization scheme, with most of the computational cost incurred in solving for the states.
2) Performance-driven methods: Several existing works estimate noise parameters under the supervision of state-estimation performance. The method in [li2019unsupervised] optimizes innovation-based objectives, using measurement residuals in the outer loop and state estimation in the inner loop. The work in [wang2023neural] proposes an adaptive disturbance estimation for a moving horizon estimator via bilevel optimization with fast derivative computation, while it requires a reference trajectory. The framework in [qadri2024learning] formulates covariance learning as constrained bilevel optimization, which requires ground-truth supervision and uses numerical gradients. Parameters can be optimized to fit the current task using these approaches. However, without probabilistic modeling, the resulting parameters may deviate from the true underlying values.
@@ III Preliminary on Robotic Lie Groups
This section presents the preliminaries on matrix Lie groups used in the derivation of this work, following [sola2018micro].
@@
Let $\mathcal{M}$ be the smooth manifold of a Lie group. For any $X_{1},X_{2}\in\mathcal{M}$ and tangent vector $\xi\in\mathbb{R}^{\dim\mathcal{M}}$, we define the plus and minus operators, $X_{1}\boxplus\xi$ and $X_{1}\boxminus X_{2}$, which map between the manifold and its tangent space. These operators admit left and right versions; the choice is application-dependent and left implicit. For instance, in the left case, $X_{1}\boxplus\xi=\exp(\xi){X}_{1},X_{1}\boxminus X_{2}=\log(X_{1}X_{2}^{-1}).$ On product manifolds, the operators act componentwise [sola2018micro]. The uncertainty of $X\in\mathcal{M}$ is modeled in the tangent space by a Gaussian distribution [barfoot2014associating]. Let $\xi\sim\mathcal{N}(0,\Sigma)$, the induced distribution is
@@ IV Problem Statement
where $X_{k}\in\mathcal{X}$ denotes the robot’s state at time step $k$ on the Lie group $\mathcal{X}$, $u_{k}\in\mathcal{U}$ is the control input, and $w_{k}$ is the process noise. Conventionally, introspective sensor measurements (e.g., from IMUs) are used as inputs to the motion model through preintegration, which conflates uncertainty in measurements with the process noise $w_{k}$ [burnett2024imu].
The motion of a robot can be captured using extrospective sensors. In this paper, we categorize their measurements into two types [dellaert2017factor]. Unary measurements, including those from GPS receivers, correspond to a single motion state and provide information in a fixed reference frame, whereas binary measurements, including LiDAR or RGB-D point clouds from iterative closest point (ICP), relate two consecutive motion states. At this stage, we do not constrain the sensor types to specific implementations to maintain theoretical completeness. The probabilistic models for these two measurement classes are
@@
As discussed in Section I, accurate knowledge of process and measurement noise covariances is crucial for odometry but remains difficult to determine in practice. Consequently, we treat these covariances as unknown quantities and introduce a parameterization via a variable $\theta$. This parameterization may take different forms, from simple vectorization of matrix entries (with symmetry, positive definiteness, and other constraints enforced) to more expressive representations based on neural networks. While some prior works [qadri2024learning, liu2025debiasing] leverage ground-truth robot poses as supervisory signals to learn the noise parameters, in this paper, we propose a general approach. We introduce a new class of measurements—termed supervisory measurements—that directly inform the robot’s relative poses over extended trajectories. These measurements provide supervision for estimating $\theta$. Formally, the supervisory measurement model is
@@ V Methodology
The assumption is standard in SLAM literature [khosoussi2025joint, zheng2024fast], which enables tractable theoretical analysis while remaining practically justifiable for most real-world applications.
@@ V-B State Filter
where $F_{k-1}=\bigl[\begin{smallmatrix}\Phi_{k-1}&\bm{0}\\ \bm{0}&I\end{smallmatrix}\bigr]$, $Q_{0,k}=\bigl[\begin{smallmatrix}Q_{0}&\bm{0}\\ \bm{0}&\bm{0}\end{smallmatrix}\bigr]$, with $\Phi_{k-1}\triangleq\bigl[\begin{smallmatrix}\bm{0}&I\\ \bm{0}&\phi_{k-1}\end{smallmatrix}\bigr]$, $Q_{0}\triangleq\bigl[\begin{smallmatrix}\bm{0}&\bm{0}\\ \bm{0}&Q\end{smallmatrix}\bigr]$. Note that $\phi_{k-1}$ follows [barrau2016invariant] and owing to the log-linear property of group-affine dynamics, $\phi_{k-1}$ is independent of ${X}_{k-1}$.
@@
By substituting the Gaussian predictive distribution from Subsection V-B into (9) and discarding additive constants, the odometry loss takes the closed-form expression [sarkka2023bayesian]
@@
Evaluating the gradient of $\ell^{o}(\theta)$ with respect to $\theta$ requires the derivatives of $S_{k}$ and $r_{k}$, with respect to $\theta$, which in turn necessitates a sensitivity analysis of the state filter. Following [sarkka2023bayesian], applying a term-wise differential to the state filter (11)–(17), which we refer to as the derivative filter.
@@ V-D Methodology Summary and More Discussions
The overall execution of the noise–covariance estimation procedure is summarized in Algorithm 1. At the upper level, the problem is posed as a constrained optimization, where the operation $\mathrm{Projection}_{\Theta}$ projects the estimated parameters onto the feasible set $\Theta$. To improve numerical stability, feasibility constraints can incorporate bounds on the eigenvalues or the condition number of the covariance matrix [qadri2024learning]. In this work, the step size $\eta^{(i)}$ is obtained using the Armijo line search rule. More generally, once the objective function and its gradient are available, a wide range of gradient-based optimization algorithms can be employed.
@@
Multiple parameterizations are possible for noise covariance matrices. A widely adopted approach is Cholesky decomposition [parellier2023speeding], which inherently guarantees positive definiteness, while the diagonal form [qadri2024learning] is a special case. Our framework is agnostic to the specific choice of parameterization and can accommodate more expressive representations, including differentiable neural network layers [wang2023neural].
@@
The proposed method is not limited to the InEKF. For a general EKF, the required derivatives can be obtained via sensitivity equations. We focus on the InEKF and affine group systems in this work, as they are prevalent in practice and possess favorable linearization properties [barrau2016invariant] that simplify the derivative–filter design.
@@
Finally, we briefly relate our approach to prior work. Noise covariance estimation is often posed as alternating optimization over states and covariances, using either a shared or separate objective at two levels. We start with the covariance-level objective $L(\theta)$, which is the log-likelihood of the parameters of interest conditioned on all measurements, as derived from a Bayesian framework. Serval prior works [qadri2024learning, liu2025debiasing, wang2023neural] tune hyperparameters, including noise covariances, by minimizing state-estimation losses against ground-truth poses, $L(\hat{\bm{X}}(\theta))$. Although effective in reducing trajectory error, the resulting covariances may not reflect the true noise statistics well. To remove reliance on ground truth, [hu2017introspective] introduced the approximate posterior error, which in the EKF reduces to the trace of the estimation error covariance, i.e., $L(\hat{\bm{P}}(\theta))$. Innovation-based methods learn covariances from measurement residuals, leading to $L(\hat{\bm{X}}(\theta))$. The traditional energy minimization approach [tsyganova2017svd, sarkka2023bayesian] corresponds to the loss $\ell^{o}(\theta)$, whereas our formulation incorporates the prevalent supervisory measurements in SLAM as an additional source of information to reinforce estimation performance. More recently, [khosoussi2025joint] proposed joint MAP estimation of states and noise covariances with closed-form covariance updates, but this framework is incompatible with filter-based state estimation and entails substantial cost due to repeated full-state optimization refinement.
@@ VI-A Baselines
    Approximate posterior error (APE) [hu2017efficient].
@@
    Mean squared error of state estimate (MSE) [qadri2024learning].
@@
    Average innovartion (Innov.) [li2019unsupervised].
@@
The implementations of gradient descent optimization in [qadri2024learning, li2019unsupervised] follow the original papers, which rely on numerical gradients. The approach of [hu2017efficient] uses Bayesian optimization for general hyperparameters, whereas we focus on a differentiable noise parameter and apply gradient descent.
@@ VI-B Simulation
We first evaluate the algorithm in simulation. The state $X_{k}\in SE_{2}(3)$ as in (1) is estimated using a left-invariant extended Kalman filter (LInEKF). The system kinematics is given by the IMU, where in LInEKF the process noise covariance is $\mathrm{blkdiag}(Q_{g},\bm{0},Q_{a})$ [hartley2020contact], with gyroscope noise $n_{g}\sim\mathcal{N}(\bm{0},Q_{g})$ and accelerometer noise $n_{a}\sim\mathcal{N}(\bm{0},Q_{a})$. The unary measurements given by GPS and binary measurements given by visual odometry (VO) are specified as
@@
A Monte Carlo experiment is conducted to evaluate the impact of supervisory measurements by comparing $\ell^{o}$ and $\mathcal{L}$ as loss functions. Measurement noises are parameterized by a base level $\beta$ and a scaling factor $\alpha$, with $\theta_{j}=(\alpha\beta_{j})^{2}$, where $\beta_{j}$ is sampled from $1,2]\times 10^{-4}$ for $Q$, $2,10]\times 10^{-3}$ for $\Sigma^{u}$, and $1,6]\times 10^{-4}$ for $\Sigma^{b}$. All algorithms are initialized identically and run for 20 iterations under the same optimizer settings. The average 2-Wasserstein error [khosoussi2025joint] is also reported, measuring the discrepancy between estimated and true covariances, with smaller values indicating more accurate estimation. As shown in Figure.5, $\mathcal{L}$ yields consistently lower MSE than $\ell^{o}$ across noise levels and, under large noise, achieves smaller Wasserstein error, indicating more accurate covariance estimation. This demonstrates the additional benefits of incorporating supervisory measurements.
@@ VI-C Dataset Experiment
The garage dataset [carlone2015initialization] provides real-world binary measurements and loop closures for evaluation. It features calibration segments with abundant loops and validation sections with open-loop trajectories. Lacking unary measurements, we use only IMU and binary data. Following convention, we treat the outlier-free estimation using all measurements as groundtruth. Noisy IMU measurements are then generated from this trajectory using known noise parameters, enabling convenient evaluation of the algorithm. As outlier rejection is not the focus of this work, we replace the false positive loop closures with measurements assigned a noise standard deviation of $0.05$. In the loop- case, 136 loop pairs are used, while in the loop+ case, 234 pairs are included.
@@ VII Conclusion and Future Work
Several aspects of the framework warrant further exploration. We plan to apply the algorithm to real-world SLAM systems such as [zheng2024fast]. Besides, we will explore deployment-oriented parameterizations, for instance, differentiable neural embeddings that adapt measurement noises to context. Moreover, the framework’s flexibility also enables tuning of broader hyperparameters, which remains to be investigated.